# Superconducting resonator parametric amplifiers with intrinsic separation of pump and signal tones

### v2

@@ III.2 Optimised Qc measurements
To verify this reasoning, we swapped the resonator of the previous section, which had a coupling length of $10\,\mathrm{\mu m}$, with a different resonator which had a coupling length of $300\,\mathrm{\mu m}$. This half-wave resonator had a quality factor of $Q_{\mathrm{r}}\sim 2000$ (when measured using a weak sweep tone). The top subfigure of figure 6 shows the resonance sweep with the same power as the operating point using the pump tone as the sweeping tone; the bottom subfigure shows the gain profile of the signal tone with the pump tone fixed at the operating point. The operating point was given by a pump power of $-12.50\,\mathrm{dBm}$ measured just outside of the cryostat, and pump frequency of $6.4726\,\mathrm{GHz}$. Far from resonance, the full power of the pump was reflected, giving a reflection coefficient of $-26.5\,\mathrm{dB}$ overall; at the operating point, where most of the pump tone was transmitted, a decreased reflection coefficient of $-38.5\,\mathrm{dB}$ was obtained instead. The pump tone in the reflection direction was reduced by $12\,\mathrm{dB}$ due to its placement near to the bottom of the resonance in reflection mode. In this way, we successfully increased the degree of pump separation by ensuring that $Q_{\mathrm{r}}$ was limited by $Q_{\mathrm{c}}$.
@@
In order to investigate the stability of the amplifier over long timescales, we fixed the signal frequency at $6.3717\,\mathrm{GHz}$, where the gain was close to $12\,\mathrm{dB}$, and monitored the gain continuously as a function of time. The measured drift was approximately $0.15\,\mathrm{dB}$ over a period of an hour. Based on our experience with resonator amplifiers, this remarkable stability can likely be further improved by reducing the power required to achieve bifurcation, which can be achieved by using more highly resistive films or by reducing the cross-sectional area of the central conductor of the CPW.
@@
Figure 9: Measurement of signal amplification and pump separation for the optimised resonator amplifier at $3.2\,\mathrm{K}$. This temperature is maintained using a pulse tuber cooler. The operating point is given by pump power of $-8.50\,\mathrm{dBm}$ measured just outside of the cryostat, and pump frequency of $6.4502\,\mathrm{GHz}$. Top subfigure: resonance response of the reflection mode where the sweep power is equal to the pump power of the operating point. Bottom subfigure: gain profile of the signal tone as a function of frequency. The pump tone is fixed in power and frequency at the operating point, and it is represented by the spike in the bottom subfigure. The peak gain of the bottom subfigure is $11\,\mathrm{dB}$. In the top subfigure, far from resonance, a reflection coefficient of $-28\,\mathrm{dB}$ is measured; at the operating point of the pump tone, a decreased reflection coefficient of $-38.5\,\mathrm{dB}$ is measured instead. Thus $10.5\,\mathrm{dB}$ of pump separation has been achieved at this operating point.
@@
Our medium-term goal is to fabricate thin-film amplifier chains that do not required HEMTs at all. We believe that this would open the door to large-format arrays that could be produced and cooled inexpensively; for example to enable the realisation of quantum-noise limited phased imaging arrays. To give an indication of whether multistage parametric amplifiers would be feasible, we operated the resonator amplifier described above at a physical temperature of $3.2\,\mathrm{K}$, which is readily achievable using a pulse tuber cooler alone. The top subfigure of figure 9 shows the resonance sweep with the same power as the operating point using the pump tone as the sweeping tone; the bottom subfigure shows the amplification profile of the signal tone with the pump tone fixed at the operating point. The operating point was achieved using a pump power of $-8.50\,\mathrm{dBm}$ measured just outside of the cryostat, and a pump frequency of $6.4502\,\mathrm{GHz}$. Under this operating condition, a peak signal gain of $11\,\mathrm{dB}$ was achieved, as shown in the lower figure. Far from resonance, the full power of the pump was reflected, giving a reflection coefficient of $-28\,\mathrm{dB}$; at the operating point, most of the pump tone was transmitted, a decreased reflection coefficient of $-38.5\,\mathrm{dB}$ was obtained instead. The pump tone in the reflection direction was reduced by $10.5\,\mathrm{dB}$ due to its placement near the resonance depth of the reflection mode. Thus both high gain and a high degree of pump separation can also be achieved at temperatures maintained by a pulse tuber cooler. Overall, this measurement demonstrated convincingly that a low-temperature amplifier, $0.1\,\mathrm{K}$, could also be operated easily, using the same set-up and method, at $3.2\,\mathrm{K}$.
@@ IV Discussion and conclusion
We have designed, realised, and investigated a scheme for configuring superconducting resonator parametric amplifiers in a way that achieves a strong separation between the pump and signal tones. Guided by our previous theoretical work on resonator parametric amplifiers [20], our scheme exploits the intrinsic asymmetry between the S-parameters of the pump and signal tones when a half-wavelength resonator is used in reflection. By utilising the reflection mode of a superconducting two-port resonator amplifier as the output through the use of a cryogenic circulator, our scheme enables strong amplification of the signal tone and strong reduction in the power of the pump tone. Compared with our previous interference-based approach, our new circulator-based scheme does not require the delicate matching and monitoring of two, long interfering paths. The new approach has turned out to be effective, rugged, and significantly reduces complexity and cost. It eliminates a source of instability caused by phase-to-amplitude conversion in the previously used interference method. Measurements on a proof-of-concept NbN amplifier has shown that the pump separation scheme works well, but the degree of separation is limited when the device is not coupling-Q limited, and so has a shallow resonance. We predicted that the degree of pump separation could be increased by increasing the lengths of the re-entrant transmission lines used to realise the coupling capacitors in our half-wavelength resonators. Our measurements on a device with longer re-entrant coupling lines confirmed this expectation. The optimised amplifier, measured at $0.1\,\mathrm{K}$, had peak signal gain of $15\,\mathrm{dB}$ whilst achieving a pump separation of $12\,\mathrm{dB}$. The amplifier was stable during continuous operation, having a gain drift of $0.15\,\mathrm{dB}$ over an hour, with no other stabilising methods employed. The same amplifier was also operated at $3.2\,\mathrm{K}$ and achieved a peak signal gain of $11\,\mathrm{dB}$ whilst having a pump separation of $10.5\,\mathrm{dB}$. These results, at both $0.1\,\mathrm{K}$ as well as $3.2\,\mathrm{K}$, compare favourably with the previously used pump cancellation method based on destructive interference. We predict that even higher degrees of pump separation should be possible by using resonators with stronger coupling. A pump separation factor of $>10\,\mathrm{dB}$ is already a highly significant result as it is sufficient to prevent the saturation of the second-stage HEMTs in typical readout systems. The simplicity of the method, together with the demonstration of good performance at $3.2\,\mathrm{K}$, opens the door to more complicated multi-channel integrated systems for quantum sensing and computing; for example of the kind that are needed for a class of inward looking quantum-noise limited phased arrays [16]. Multistage amplifiers that eliminate the need for a HEMT amplifier altogether are starting to look feasible.

### v3

@@ III.3 Optimised Qc measurements
To verify this reasoning, we swapped the resonator of the previous section, which had a coupling length of $10\,\mathrm{\mu m}$, with a different resonator which had a coupling length of $300\,\mathrm{\mu m}$. This half-wave resonator had a quality factor of $Q_{\mathrm{r}}\sim 2000$ (when measured using a weak sweep tone). The top subfigure of figure 6 shows frequency sweep of the resonance using the pump tone as the sweeping tone. The pump tone has the same power as the amplifier operating point and is the only tone present in this resonance sweep. The bottom subfigure shows the gain profile of the signal tone with the pump tone fixed at the operating point. The operating point was given by a pump power of $-12.50\,\mathrm{dBm}$ measured just outside of the cryostat, and pump frequency of $6.4726\,\mathrm{GHz}$. Far from resonance, the full power of the pump was reflected, giving a reflection coefficient of $-26.5\,\mathrm{dB}$ overall; at the operating point, where most of the pump tone was transmitted, a decreased reflection coefficient of $-38.5\,\mathrm{dB}$ was obtained instead. The pump tone in the reflection direction was reduced by $12\,\mathrm{dB}$ due to its placement near to the bottom of the resonance in reflection mode. In this way, we successfully increased the degree of pump separation by ensuring that $Q_{\mathrm{r}}$ was limited by $Q_{\mathrm{c}}$.
@@
In order to investigate the characteristic stability of the amplifier over hour-long timescales, we fixed the signal frequency at $6.3717\,\mathrm{GHz}$, where the gain was close to $12\,\mathrm{dB}$, and monitored the gain continuously as a function of time. The measured drift was approximately $0.15\,\mathrm{dB}$ over a period of an hour. Based on our experience with resonator amplifiers, this remarkable stability can likely be further improved by reducing the power required to achieve bifurcation, which can be achieved by using more highly resistive films or by reducing the cross-sectional area of the central conductor of the CPW.
@@
Figure 9: Measurement of signal amplification and pump separation for the optimised resonator amplifier at $3.2\,\mathrm{K}$. This temperature is maintained using a pulse tube cooler. The operating point is given by pump power of $-8.50\,\mathrm{dBm}$ measured just outside of the cryostat, and pump frequency of $6.4502\,\mathrm{GHz}$. Top subfigure: resonance response of the reflection mode where the sweep power is equal to the pump power of the operating point. Bottom subfigure: gain profile of the signal tone as a function of frequency. The pump tone is fixed in power and frequency at the operating point, and it is represented by the spike in the bottom subfigure. The peak gain of the bottom subfigure is $11\,\mathrm{dB}$. In the top subfigure, far from resonance, a reflection coefficient of $-28\,\mathrm{dB}$ is measured; at the operating point of the pump tone, a decreased reflection coefficient of $-38.5\,\mathrm{dB}$ is measured instead. Thus $10.5\,\mathrm{dB}$ of pump separation has been achieved at this operating point.
@@
Our medium-term goal is to fabricate thin-film amplifier chains that do not require HEMTs at all. We believe that this would open the door to large-format arrays that could be produced and cooled inexpensively; for example to enable the realisation of quantum-noise limited phased imaging arrays. Developing quantum limited amplifiers that reliably operate at pulse tube cooler temperatures and above is an important theme in the ongoing research of parametric amplifiers [30, 31]. To give an indication of whether multistage parametric amplifiers would be feasible, we operated the resonator amplifier described above at a physical temperature of $3.2\,\mathrm{K}$, which is readily achievable using a pulse tube cooler alone. The top subfigure of figure 9 shows frequency sweep of the resonance using the pump tone as the sweeping tone. The pump tone has the same power as the amplifier operating point and is the only tone present in this resonance sweep. The bottom subfigure shows the amplification profile of the signal tone with the pump tone fixed at the operating point. The operating point was achieved using a pump power of $-8.50\,\mathrm{dBm}$ measured just outside of the cryostat, and a pump frequency of $6.4502\,\mathrm{GHz}$. With these operating conditions, a peak signal gain of $11\,\mathrm{dB}$ was achieved, as shown in the lower figure. Far from resonance, the full power of the pump was reflected, giving a reflection coefficient of $-28\,\mathrm{dB}$; at the operating point, most of the pump tone was transmitted, a decreased reflection coefficient of $-38.5\,\mathrm{dB}$ was obtained instead. The pump tone in the reflection direction was reduced by $10.5\,\mathrm{dB}$ due to its placement near the resonance depth of the reflection mode. Thus both high gain and a high degree of pump separation can also be achieved at temperatures maintained by a pulse tube cooler. Overall, this measurement demonstrated convincingly that a low-temperature amplifier, $0.1\,\mathrm{K}$, could also be operated easily, using the same set-up and method, at $3.2\,\mathrm{K}$.
@@ IV Discussion and conclusion
We have designed, realised, and investigated a scheme for configuring superconducting resonator parametric amplifiers in a way that achieves a strong separation between the pump and signal tones. Guided by our previous theoretical work on resonator parametric amplifiers [22], our scheme exploits the intrinsic asymmetry between the S-parameters of the pump and signal tones when a half-wavelength resonator is used in reflection. By utilising the reflection mode of a superconducting two-port resonator amplifier as the output through the use of a cryogenic circulator, our scheme enables strong amplification of the signal tone and strong reduction in the power of the pump tone. Compared with our previous interference-based approach, our new circulator-based scheme does not require the delicate matching and monitoring of two, long interfering paths. The new approach has turned out to be effective, rugged, and significantly reduces complexity and cost. It eliminates a source of instability caused by phase-to-amplitude conversion in the previously used interference method. Measurements on a proof-of-concept NbN amplifier has shown that the pump separation scheme works well, but the degree of separation is limited when the device is not coupling-Q limited, and so has a shallow resonance. We predicted that the degree of pump separation could be increased by increasing the lengths of the re-entrant transmission lines used to realise the coupling capacitors in our half-wavelength resonators. Our measurements on a device with longer re-entrant coupling lines confirmed this expectation. The optimised amplifier, measured at $0.1\,\mathrm{K}$, had peak signal gain of $15\,\mathrm{dB}$ whilst achieving a pump separation of $12\,\mathrm{dB}$. The amplifier was stable during continuous operation, having a gain drift of $0.15\,\mathrm{dB}$ over an hour, with no other stabilising methods employed. The same amplifier was also operated at $3.2\,\mathrm{K}$ and achieved a peak signal gain of $11\,\mathrm{dB}$ whilst having a pump separation of $10.5\,\mathrm{dB}$. These results, at both $0.1\,\mathrm{K}$ as well as $3.2\,\mathrm{K}$, compare favourably with the previously used pump cancellation method based on destructive interference. We predict that even higher degrees of pump separation should be possible by using resonators with stronger coupling. A pump separation factor of $>10\,\mathrm{dB}$ is already a highly significant result as it is sufficient to prevent the saturation of the second-stage HEMTs in typical readout systems. The simplicity of the method, together with the demonstration of good performance at $3.2\,\mathrm{K}$, opens the door to more complicated multi-channel integrated systems for quantum sensing and computing; for example of the kind that are needed for a class of inward looking quantum-noise limited phased arrays [20]. Multistage amplifiers that eliminate the need for a HEMT amplifier altogether are starting to look feasible. Future research should be carried out to build on the core idea of exploiting the asymmetry between the pump and the signal tones whilst exploring more sophisticated configurations of resonators and couplers. For example, improving the bandwidth of kinetic inductance resonator amplifiers through the use of impedance transformers, a technique previously demonstrated with great success in the context of Josephson Parametric Amplifiers [32, 33], will improve the utility of these devices. Finally, measurement of the added noise from different configurations of these resonator parametric amplifiers will be important in evaluating the optimal configuration for each specific context of applications.
Acknowledgements.
The authors are grateful for funding from the UK Research and Innovation (UKRI) and the Science and Technology Facilities Council (STFC) through the Quantum Technologies for Fundamental Physics (QTFP) programme (Project Reference ST/T006307/2).